Please help solve this:16(x-1/x-8)²+8(x-1/x-8)+1=0
Answer by Alan3354(69443)

Please help solve this:16(x-1/x-8)²+8(x-1/x-8)+1=0
-------------
Assuming (x-1/x-8) means (x-1)/(x-8):
Sub y for (x-1)/(x-8)
-----
16y^2 + 8y + 1 = 0
(4y+1)^2 = 0
y = -1/4
-----------------
(x-1)/(x-8) = -1/4
x-1 = -(x-8)/4
4x-4 = -x+8
5x = 12
x = 12/5